Solve the system using substitution
3x+2y=9
x=4-y

Answer by jim_thompson5910(35256)

3x+2y=9

3(4-y)+2y=9

12-3y+2y=9

-y+12=9

-y=9-12

-y=-3

y=(-3)/(-1)

y=3
--------------------
x = 4-y

x = 4-3

x = 1


So the solutions are x = 1 and y = 3


This means that the system is consistent and independent.